OpenStudy (anonymous):

The maximum speed of the Tupolev Tu-144 airliner is 694 m/s. What is this speed in kilometers per hour?

OpenStudy (anonymous):

the easiest trick would be to multiply by each conversion one by one like so: \[\frac{ 694 m }{ 1 s }\times \frac{ 1 km }{ 1000 m }\times \frac{ 60 s }{ 1 \min }\times \frac{ 60 \min }{ 1 hour } = ?\] you can make sure you are multiplying by the right conversion factor by making sure that in each step your units are canceling (eg: notice how the meters cancel out here)